An auto's velocity increase uniformly form 7.4 m/s to 20 m/s while covering 70m in straight line. Find the acceleration and the time taken.


1
Expert's answer
2021-10-26T09:39:04-0400

We can find the acceleration of the car as follows:


"v^2=v_0^2+2ad,""a=\dfrac{v^2-v_0^2}{2d},""a=\dfrac{(20\ \dfrac{m}{s})^2-(7.4\ \dfrac{m}{s})^2}{2\times70\ m}=2.47\ \dfrac{m}{s^2}."

We can find the time taken as follows:


"v=v_0+at,""t=\dfrac{v-v_0}{a}=\dfrac{20\ \dfrac{m}{s}-7.4\ \dfrac{m}{s}}{2.47\ \dfrac{m}{s^2}}=5.1\ s."